OpenStudy (anonymous):

please help!! \[2\left| 2x-6 \right|-4 <10\]

OpenStudy (anonymous):

-7<2x - 6<7 -1<2x<13 -1/2<x<13/2

OpenStudy (anonymous):

oh great man i forgot the sign when i was doing it xD

OpenStudy (anonymous):

\[\left| x-5 \right| -14 = -12\] what about that ?

OpenStudy (anonymous):

hmm |x - 5| = 2 x = 3, or 7